OpenStudy (anonymous):

if 3 wergs equals 4 nergs and 6 nergs equal 7 clergs, then 5 wergs equal how many clergs?

OpenStudy (anonymous):

Lets say a werg is represented by the letter w, a nerg is n, and a clerg is c. The word problem tells us that 3w = 4n and 6n = 7c. So w = (4/3)n n = (7/6)c so 6 w = 6*(4/3)n = 6*(4/3)*(7/6)c use a calculator to do the math

OpenStudy (anonymous):

matt are u still there?